[ A car travels 432 km on 48 liters of petrol. How far will it travel on 20 liter of petrol? ]
Answer by TimothyLamb(4379)

---
d/v = (432/48)
d = (432/48) * v
d = (432/48) * 20
d = 180 km
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php